Assuming that a household in Happy-land consumes only apple juice and T-shirts. Last year, which was the base year, the household spent $440 on juice and $132 on T-shirts. In the base year, the juice was $2.2 a bottle and T-shirts were $11. This year, the juice is $3.3 a bottle, T-shirts are $13.2 each, and a typical household has bought 198 bottles of juice and 15 T-shirts. Is the inflation rate that you've calculated likely to be biased? Why or why not?
1
Expert's answer
2020-07-17T10:32:11-0400

the base year

"440+132=572" income

this year

"198*3.3=653.4"

"15*13.2=198"

"653.4+198=851.4" income

the inflation rate :

"((\frac{3.3}{2.2}-1)*\frac{653.4}{851.4}+(\frac{13.2}{11}-1)*\frac{198}{851.4})/(\frac{851.4}{572}-1)-1=-0.12"

We have deflation 12%

My calculation is not biased, because I considered both occasions the price increase and the consumers income increase